OpenStudy (sighn0more):

Hello! Please help me work through and solve this problem (it requires the quadratic equation): (3x + 1) ^2 = -2x

OpenStudy (sighn0more):

Do I need to expand the binomial first or does this problem require a different course?

OpenStudy (freckles):

yes

OpenStudy (sighn0more):

Okay, so that would get me to 9x^2 + 6x + 2 = -2x I think

OpenStudy (freckles):

\[(3x+1)^2=(3x)^2+2(3x)(1)+1^2 \\ (3x+1)^2=9x^2+6x+1\]

OpenStudy (freckles):

so should be: \[9x^2+6x+1=-2x \\ \text{ then add } 2x \text{ on both sides } \\ 9x^2+6x+2x+1=0 \\ 9x^2+8x+1=0\]

OpenStudy (sighn0more):

Oh yes, sorry. Error on my part! And from here should these numbers be my values for a, b, and c?

OpenStudy (freckles):

yes you can use quadratic formula if you want a=9 b=8 c=1 in the end you should check your answers x being positive will not work just so you know because a real number squared will result in a positive number (or zero) (3x+1)^2=-2x we need x to be negative because 0 obviously doesn't work either

OpenStudy (freckles):

and after solving these equation I can tell you you will not have any domain issues with the answers you get

OpenStudy (sighn0more):

Perfect! I ended up getting -8 +- radical 28 / 18

OpenStudy (sighn0more):

I guess I would need to simplify this further...

OpenStudy (freckles):

\[x=\frac{-8 \pm \sqrt{28}}{18} \\ \text{ note : } \sqrt{28}=\sqrt{4 \cdot 7} =\sqrt{4} \sqrt{7}=2 \sqrt{7} \\ \] yes your solution can be simplified see the note as a hint

OpenStudy (sighn0more):

Oh, so my final answer would be x = -4/9 +- radical 7 / 9

OpenStudy (freckles):

\[x=\frac{-8 \pm 2 \sqrt{7}}{18} \\ \text{ divide \top and bottom by 2 } \\ x=\frac{\frac{-8}{2} \pm \frac{2}{2} \sqrt{7}}{\frac{18}{2}} \\ x=\frac{-4 \pm \sqrt{7}}{9} \\ \text{ yes this could be written as } x=\frac{-4}{9} \pm \frac{\sqrt{7}}{9}\]

OpenStudy (sighn0more):

Thank you!!! This helped a bunch.

OpenStudy (freckles):

cool stuff! :)